if you know the names of the remaining 8 students in the spelling bee, what is the probability of randomly selecting an order and getting the order that is used in the spelling bee?
Answer by Theo(13342)

there are 40320 possible arrangements.

that would be 8!

the probability of picking the correct order randomly would be 1 / 40320 = .0000248015873.

in order words, the probability of picking the correct order on a random basis, such as picking the names out of a hat, is not very good.

for example:

suppose there were only 3 reamining candidate.

3! = 3*2 = 6

there would be 6 possible arrangements.

if the candidates were abc, then the 6 possible arrangements would be:

abc
abc
bac
bca
cab
cba

for 4 candidates, multiply 6 * 4 = 24
then 5 * 24 = 120
then 6*120 = 720
then 7*720 = 5040
then 8 * 5040 = 40320

1 / 6 is manageable.
1 / 24 less.
1 / 8 becomes pretty impossible.